OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9 PLEASE I NEED THIS DONE FAST! MEDAL AND FAN!

OpenStudy (igreen):

4x - 5y = 9 Subtract 4x to both sides: -5y = -4x + 9 Divide -5 to all terms: y = -4/5x + 9

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (igreen):

No problem.